# Optimal Learners for Realizable Regression: PAC Learning and Online Learning

### v1

@@ B.1 Partial Concept Classes and A Naive Approach that Fails
Predicting normal-$\star$ in [AHHM22] implies a mistake for the setting of partial concept classes. However, in our regression setting, normal-$\star$ is interpreted differently and corresponds to loss at most $\gamma$ which is desirable. In particular, the hard instance for proper learners in the partial concepts paper (see Proposition 4 in [AHHM22]) is good in settings where predicting normal-$\star$ does not count as a mistake, as in our regression case.
@@ B.2 Main Result
$\Omega\left(\frac{\mathbb{D}_{\gamma}^{\mathrm{G}}(\mathcal{H})+\log(1/\delta)}{\varepsilon}\right)\leq\mathcal{M}_{\mathrm{ERM}}(\mathcal{H};\varepsilon,\delta)\leq O\left(\frac{\mathbb{D}_{\gamma}^{\mathrm{G}}(\mathcal{H})\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon}\right)\,.$
@@ Appendix C $\gamma$-OIG Dimension and Learnability
In this section we identify a dimension characterizing qualitatively and quantitatively what classes of predictors $\mathcal{H}\subseteq[0,1]^{\mathcal{X}}$ are PAC learnable and we provide PAC learners that achieve (almost) optimal sample complexity. In particular, we show the following result.
@@ C.2.1 Scaled One-Inclusion Graph Algorithm and Weak Learning
First, we need to create an orientation of the whole (potentially infinite) one-inclusion graph.
We will create this orientation using the compactness theorem of first-order logic which states that a set of formulas $\Phi$ is satisfiable if and only if it is finitely satisfiable, i.e., every finite subset $\Phi^{\prime}\subseteq\Phi$ is satisfiable. Let $G^{\mathrm{OIG}}_{\mathcal{H}|_{S}}=(V_{n},E_{n})$ be the (potentially infinite) one-inclusion graph of $\mathcal{H}|_{S}$. Let $\mathcal{Z}$ be the set of pairs $z=(v,e)\in V_{n}\times E_{n}$ so that $v\in e$. Our goal is to assign binary values to each $z\in\mathcal{Z}.$ We define the following sets of formulas:
@@ C.2.3 Generalization via Sample Compression Schemes
Sample compression scheme is a classic technique for proving generalization bounds, introduced by [LW86, FW95]. These bounds proved to be useful in numerous learning settings, such as binary classification [GHST05, MY16, BHMZ20], multiclass classification [DSBDSS15, DSS14, DMY16, ${}^{+}$22], regression [HKS18, HKS19], active learning [WHEY15], density estimation [${}^{+}$20], adversarially robust learning [MHS19, MHS20, MHS21, MHS22, AHM22, AH22], learning with partial concepts [AHHM22], and showing Bayes-consistency for nearest-neighbor methods [GKN14, KSW17]. As a matter of fact, compressibility and learnability are known to be equivalent for general learning problems [DMY16]. Another remarkable result by [MY16] showed that VC classes enjoy a sample compression that is independent of the sample size.
@@
A pair of functions $(\kappa,\rho)$ is a sample compression scheme of size $\ell$ for class $\mathcal{H}$ if for any $n\in\mathbb{N}$, $h\in\mathcal{H}$ and sample $S=\{(x_{i},h(x_{i}))\}^{n}_{i=1}$, it holds for the compression function that $\kappa\left(S\right)\subseteq S$ and $|\kappa\left(S\right)|\leq\ell$, and the reconstruction function $-^\rho\left(\kappa\left(S\right)\right)=\hat{h}$ satisfies $\hat{h}(x_{i})=h(x_{i})$ for any $i\in[n]$.
@@ Appendix D $\gamma$-DS Dimension and Learnability
We believe that there must exist a modification of the approach of [${}^{+}$22] that will be helpful in settling the above conjecture.
@@
There exists $\mathcal{H}\subseteq(0,1)^{\mathcal{X}}$ for which $\mathbb{D}^{\mathrm{Nat}}_{\gamma}(\mathcal{H})=1$ for all $\gamma\in(0,1)$ but $\mathbb{D}^{\mathrm{DS}}_{\gamma}(\mathcal{H})<\infty$ for some $\gamma\in(0,1)$.
In particular, we believe that one can extend the construction of [${}^{+}$22] (which uses various tools from algebraic topology as a black-box) and obtain a hypothesis class $\mathcal{H}\subseteq[0,1]^{\mathcal{X}}$ that has $\gamma$-Natarajan dimension 1 but is not PAC learnable (it will have infinite $\gamma$-DS dimension). This construction though is not immediate and requires new ideas related to the works of [JŚ03, Osa13]
@@ Appendix F Dimension and Finite Character Property
[${}^{+}$19] gave a formal definition of the notion of “dimension” or “complexity measure”, that all previously proposed dimensions in statistical learning theory comply with. In addition to characterizing learnability, a dimension should satisfy the finite character property:
Definition 17 (Finite Character [${}^{+}$19]).
@@ Appendix G Examples for Scaled Graph Dimension
[…]
@@ Appendix H Extension to More General Loss Functions
Our results can be extended to loss functions that satisfy approximate pseudo-metric axioms (see e.g., [HKLM22, CKW08]). The main difference from metric losses is that we allow an approximate triangle inequality instead of a strict inequality. Many natural loss functions are captured by this definition, such as the well-studied $\ell_{p}$ losses for the regression setting. Abstractly, in this context, the label space is an abstract non-empty set $\mathcal{Y}$, equipped with a general loss function $\ell:\mathcal{Y}^{2}\to\mathbb{R}_{\geq 0}$ satisfying the following property.

### v3

@@ B.1 Partial Concept Classes and A Naive Approach that Fails
Predicting $\star$ in [AHHM22] implies a mistake for the setting of partial concept classes. However, in our regression setting, $\star$ is interpreted differently and corresponds to loss at most $\gamma$ which is desirable. In particular, the hard instance for proper learners in the partial concepts paper (see Proposition 4 in [AHHM22]) is good in settings where predicting $\star$ does not count as a mistake, as in our regression case.
@@ B.2 Main Result
$\Omega\left(\frac{\mathbb{D}_{\gamma}^{\mathrm{G}}(\mathcal{H})+\log(1/\delta)}{\varepsilon}\right)\leq\mathcal{M}_{\mathrm{ERM}}(\mathcal{H};\varepsilon,\delta,\gamma)\leq O\left(\frac{\mathbb{D}_{\gamma}^{\mathrm{G}}(\mathcal{H})\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon}\right)\,.$
@@ Appendix C $\gamma$-OIG Dimension and Learnability
In this section, we identify a dimension characterizing qualitatively and quantitatively what classes of predictors $\mathcal{H}\subseteq[0,1]^{\mathcal{X}}$ are PAC learnable and we provide PAC learners that achieve (almost) optimal sample complexity. In particular, we show the following result.
@@ C.2.1 Scaled One-Inclusion Graph Algorithm and Weak Learning
First, we need to create an orientation of the whole (potentially infinite) one-inclusion graph. We will create this orientation using the compactness theorem of first-order logic which states that a set of formulas $\Phi$ is satisfiable if and only if it is finitely satisfiable, i.e., every finite subset $\Phi^{\prime}\subseteq\Phi$ is satisfiable. Let $G^{\mathrm{OIG}}_{\mathcal{H}|_{S}}=(V_{n},E_{n})$ be the (potentially infinite) one-inclusion graph of $\mathcal{H}|_{S}$. Let $\mathcal{Z}$ be the set of pairs $z=(v,e)\in V_{n}\times E_{n}$ so that $v\in e$. Our goal is to assign binary values to each $z\in\mathcal{Z}.$ We define the following sets of formulas:
@@ C.2.2 Boosting Real-Valued Functions
  1. 1.
    Find a $(\gamma,\beta)$-weak learner $h_{t}$ with respect to $(x_{i},y_{i})\sim\mathcal{P}_{t}$, using a subset $S_{t}\subseteq S$.
  2. 2.
    For $i=1,\ldots,m$:
    1. (a)
      Set $w^{(t)}_{i}=1-2\mathbb{I}\mathopen{}\big{[}\ell(h_{t}(x_{i}),y_{i})>\gamma\big{]}.$
    2. (b)
      Set $\alpha_{t}=\frac{1}{2}\log\mathopen{}\left(\frac{\mathopen{}\left(1-\gamma\right)\sum^{n}_{i=1}\mathcal{P}_{t}\mathopen{}\left(x_{i},y_{i}\right)\mathbb{I}\mathopen{}\left[w_{i}^{(t)}=1\right]}{\mathopen{}\left(1+\gamma\right)\sum^{n}_{i=1}\mathcal{P}_{t}\mathopen{}\left(x_{i},y_{i}\right)\mathbb{I}\mathopen{}\left[w_{i}^{(t)}=-1\right]}\right).$
    3. (c)
      • If $\alpha_{t}=\infty$: return $T$ copies of $h_{t}$, $\mathopen{}\left(\alpha_{1}=1,\ldots,\alpha_{T}=1\right)$, and $S_{t}$.
      • Else: $P_{t+1}(x_{i},y_{i})=P_{t}(x_{i},y_{i})\exp\mathopen{}\left(-\alpha_{t}w_{i}^{t}\right)/Z_{t}$, where $Z_{t}=\sum^{n}_{j=1}\mathcal{P}_{t}\mathopen{}\left(x_{j},y_{j}\right)\exp\mathopen{}\left(-\alpha_{t}w_{j}^{t}\right)$.
@@ C.2.3 Generalization via Sample Compression Schemes
Sample compression scheme is a classic technique for proving generalization bounds, introduced by [LW86, FW95]. These bounds proved to be useful in numerous learning settings, such as binary classification [GHST05, MY16, BHMZ20], multiclass classification [DSBDSS15, DSS14, DMY16, BCD+22], regression [HKS18, HKS19], active learning [WHEY15], density estimation [ABDH+20], adversarially robust learning [MHS19, MHS20, MHS21, MHS22, AHM22, AH22], learning with partial concepts [AHHM22], and showing Bayes-consistency for nearest-neighbor methods [GKN14, KSW17]. As a matter of fact, compressibility and learnability are known to be equivalent for general learning problems [DMY16]. Another remarkable result by [MY16] showed that VC classes enjoy a sample compression that is independent of the sample size.
@@
A pair of functions $(\kappa,\rho)$ is a sample compression scheme of size $\ell$ for class $\mathcal{H}$ if for any $n\in\mathbb{N}$, $h\in\mathcal{H}$ and sample $S=\{(x_{i},h(x_{i}))\}^{n}_{i=1}$, it holds for the compression function that $\kappa\left(S\right)\subseteq S$ and $|\kappa\left(S\right)|\leq\ell$, and the reconstruction function $\rho\left(\kappa\left(S\right)\right)=\hat{h}$ satisfies $\hat{h}(x_{i})=h(x_{i})$ for any $i\in[n]$.
@@ Appendix D $\gamma$-DS Dimension and Learnability
We believe that there must exist a modification of the approach of [BCD+22] that will be helpful in settling the above conjecture.
@@
There exists $\mathcal{H}\subseteq(0,1)^{\mathcal{X}}$ for which $\mathbb{D}^{\mathrm{Nat}}_{\gamma}(\mathcal{H})=1$ for all $\gamma\in(0,1)$ but $\mathbb{D}^{\mathrm{DS}}_{\gamma}(\mathcal{H})=\infty$ for some $\gamma\in(0,1)$.
In particular, we believe that one can extend the construction of [BCD+22] (which uses various tools from algebraic topology as a black-box) and obtain a hypothesis class $\mathcal{H}\subseteq[0,1]^{\mathcal{X}}$ that has $\gamma$-Natarajan dimension 1 but is not PAC learnable (it will have infinite $\gamma$-DS dimension). This construction though is not immediate and requires new ideas related to the works of [JŚ03, Osa13]
@@ Appendix F Dimension and Finite Character Property
[BDHM+19] gave a formal definition of the notion of “dimension” or “complexity measure”, that all previously proposed dimensions in statistical learning theory comply with. In addition to characterizing learnability, a dimension should satisfy the finite character property:
Definition 17 (Finite Character [BDHM+19]).
@@ Appendix G Examples for Scaled Graph Dimension
    The second algorithm either returns the all-zero function or identifies the correct hypothesis if the sample contains a non-zero label. This is a good ERM learner $\mathcal{A}_{\text{good}}^{\text{ERM}}$ with sample complexity $m(\varepsilon,\delta)=\frac{1}{\varepsilon}\log\left(\frac{1}{\delta}\right).$
@@ Appendix H Extension to More General Loss Functions
Our results can be extended to loss functions that satisfy approximate pseudo-metric axioms (see e.g., [HKLM22, CKW08]). The main difference from metric losses is that we allow an approximate triangle inequality instead of a strict inequality. Many natural loss functions are captured by this definition, such as the well-studied $\ell_{p}$ losses for the regression setting. Abstractly, in this context, the label space [^12] We would like to mention that, in general, we do not require that the label space is bounded. In contrast, we have to assume that the loss function takes values in a bounded space. This is actually necessary since having an unbounded loss in the regression task would potentially make the learning task impossible. For instance, having some fixed accuracy goal, one could construct a learning instance (distribution over labeled examples) that would make estimation with that level of accuracy trivially impossible. is an abstract non-empty set $\mathcal{Y}$, equipped with a general loss function $\ell:\mathcal{Y}^{2}\to\mathbb{R}_{\geq 0}$ satisfying the following property.